Nyaanyaa:

A vertical pole is supported by two ropes staked to the ground on opposite sides of the pole one rope is 8 meters long and the other is 7 meters the distance between the stakes is 6 meters and the height of the pole is meters

Vocaloid:

It's a fill in the blank question, asking for the height of the pole in meters Answer choices: a) 3.30 b) 3.78 c) 6.17 d) 6.78 I'll let someone else have a chance first 'cause I'm a little busy rn

karissafrazier:

i am legit so bad at math i have to bother @nishio about helping me i'm just here cause i am

mhanifa:

Best way is to visualize the problem. 

mhanifa:

You can set two equations based on Pythagorean and solve for x

surjithayer:



surjithayer:

\[x^2=8^2-y^2 ...(1)\] \[x^2=7^2-(6-y)^2\] \[64-y^2=49-(6-y)^2\] \[64-49=y^2-(6-y)^2\] \[15=y^2-(36-12y+y^2)\] \[15=y^2-36+12y-y^2\] 12y=15+36=51 \[y=\frac{ 51 }{ 12 }=\frac{ 17 }{ 4 }\] now find x using (1)